OpenStudy (anonymous):

Write a program to compute the 394852th digit of pi.

hero (hero):

Okay.....good luck with that...

OpenStudy (anonymous):

use\[\pi=\sum_{i=0}^{\infty}{\frac{1}{16^i}(\frac{4}{8i+1}-\frac{2}{8i+4}-\frac{1}{8i+5}-\frac{1}{8i+6})}\]

OpenStudy (anonymous):

x-intercept | 0 theta-intercept | 0 slope | 1/(394852 pi)~~8.0615x10^-7

OpenStudy (anonymous):

what's a theta intercept

OpenStudy (anonymous):

oh